OpenStudy (anonymous):

 A dehydrated patient needs a 7.9% saline IV. Unfortunately, the hospital only has bags of 3% and 10% saline solutions. How many liters of each of these solutions should be mixed together to yield 3 liters of the desired concentration?

OpenStudy (mertsj):

x liters of 3% 3-x liters of 10$ .03x + .10(3-x)= .079(3)

OpenStudy (anonymous):

ok do I solve for x and y? to get the concentrations from the 3& and the 10% bags?

OpenStudy (mertsj):

30x+100(3-x)=237 30x+300-100x=237 -70x=-63 x=.9 liters of 3% solution 3-x = 2.1 liters of 10 % solution

OpenStudy (mertsj):

First multiply by 1000 to get rid of the decimals.

OpenStudy (anonymous):

ok , I got x = 0.9

OpenStudy (anonymous):

so that will be concentration from the 3% bag right? then I need to find the concentration from the 10% bag correct?

OpenStudy (mertsj):

Good!! That is the amount of 3% solution that is needed.

OpenStudy (mertsj):

Exactly

OpenStudy (anonymous):

ok ty

OpenStudy (mertsj):

yw

hero (hero):

my approach works, I just make mistakes

hero (hero):

x + y = 3 .03x + .1y = .079(3)